# Evolution of the Inner Accretion Flow in Swift J1727.8–1613 across Intermediate States: Insights from Broadband Spectral and Timing Analysis

### v1

@@ 3.1.1 HIMS
The clear excess at $\gtrsim 100$ keV further suggests that the models with thermal Comptonization and disk reflection alone are insufficient to explain the high-energy emission, and may originate due to the presence of a power-law tail originating in the non-thermal distribution of hot electrons in the corona. We then tried to model this power-law tail by invoking a hybrid Comptonization scenario within the compps framework, wherein the Lorentz factor ($\gamma_{\rm{min}}$) defines the transition point from a thermal to a power-law electron distribution. We allowed $\gamma_{\rm{min}}$ to vary freely while fixing $\gamma_{\rm{max}}=1000$ and keeping the electron power-law index fixed at 2. This model, referred to as M3, which assumes a hybrid electron distribution with Maxwellian and non-thermal above a certain value of $\gamma_{\rm{min}}$, provides a very good and significantly improved spectral fit in the 0.72000.7-200 keV band with $\chi^{2}$/dof=601.5/681=601.5/681 with $\Delta\chi^{2}\approx-285$ for 1 dof. The best-fit spectral parameters obtained from M3 are listed in Table 2, and the best-fit spectral data with the model components are shown in Figure 4. We find $\gamma_{\rm{min}}$ to be $\sim 1.33$, implying a significant contribution from the non-thermal distribution. The other best-fit spectral parameters obtained with model M3 agree closely with those obtained from our earlier models, assuming a purely thermal electron distribution up to $78$ keV.
3.1.2 SIMS
@@
Thcomp $\Gamma$ $2.36\pm 0.03$ $2.33^{+0.06}_{-0.03}$ $2.35\pm 0.04$
$kT_{\rm{e}}$ (keV) $998.2_{-500.4}$ $462.5_{-382.3}$ $1000_{-505.8}$
$C_{\rm{f}}$ $0.25^{+0.04}_{-0.06}$ $0.2\pm 0.1$ $0.23^{+0.06}_{-0.1}$
Kerrbb $a$  $\cdots$ $0.79^{+0.03}_{-0.13}$  $\cdots$
$i^{\circ}$  $\cdots$ $46.6^{+3.2}_{-4.7}$  $\cdots$
$\dot{M}$ ($\times 10^{18}~\rm{g/s}$)  $\cdots$ $1.52\pm 0.08$  $\cdots$
$M_{\rm{BH}}$ ($M_{\odot}$)  $\cdots$ $10.5^{+7.7}_{-3.0}$  $\cdots$
$D$ (kpc)  $\cdots$ $3.5^{+1.3}_{-0.5}$  $\cdots$
BHSPEC $\alpha=0.1$] $M_{\rm{BH}}$ ($M_{\odot}$)  $\cdots$  $\cdots$ $19.5^{+1.4}_{-6.0}$
$L/L_{\rm{Edd}}$  $\cdots$  $\cdots$ $0.19^{+0.03}_{-0.10}$
$a$  $\cdots$  $\cdots$ $0.8_{-0.08}$
$i^{\circ}$  $\cdots$  $\cdots$ $47.4^{+2.7}_{-5.0}$
$D$ (kpc)  $\cdots$  $\cdots$ $6.0_{-0.8}$
RelxillCp (refl. only) $a$ $0.98_{-0.29}$  $\cdots$  $\cdots$
$i^{\circ}$ $40^{f}$  $\cdots$  $\cdots$
$R_{\rm{in}}$ ($r_{\mathrm{isco}}$) $1.7^{+0.9}$ $1^{f}$ $1^{f}$
$q$ $3.7^{+0.6}_{-0.4}$ $6.55_{-2.99}$ $6.14^{+2.8}_{-2.2}$
$\log\xi$ $4.09^{+0.04}_{-0.11}$ $4.27^{+0.13}_{-0.25}$ $4.20^{+0.07}_{-0.12}$
@@
$K_{\rm{NICER}}$ $0.82\pm 0.02$ $0.81\pm 0.02$ $0.81\pm 0.02$
$\Delta\Gamma_{\rm{NuSTAR~B}}$ $002^{+0.006} 0.0$$03^{+0.006} 0.0$$002^{+0.006}$
@@
  • Note. M4: plabs*tbabs(thcomp*diskbb+mbknpo*relxillCp), M5: plabs*tbabs(thcomp*kerrbb+mbknpo*relxillCp), M6: plabs*tbabs(thcomp*bhspec+mbknpo*relxillCp). f – indicates fixed parameters.
@@
We then employed the more physically motivated and relativistically blurred reflection model, relxillCp, to perform broadband reflection spectroscopy, replacing the ireflect and gaussian components. Here, relxillCp is considered only as a reflection component by fixing the reflection fraction parameter at 1-1. The common parameters such as $\Gamma$ and $kT_{\rm{e}}$ were tied between the relxillCp and thcomp components. The inclination angle and the iron abundance were also kept fixed at $40^{\circ}$ and solar, respectively. Other free spectral parameters of the relxillCp model are the emissivity index ($q$), log of disk ionization ($\log\xi$), log of disk density ($\log\xi$), and normalization. Additionally, we varied the black hole spin ($a$) and inner disk radius ($R_{\rm{in}}$) freely, considering the SIMS. The resulting model plabs*tbabs(thcomp*diskbb+relxillCp) (now referred to as M4) provides a good and statistically acceptable fit with $\chi^{2}/dof=506.5/521$. This further suggests that a model consisting of a single Comptonizing region, its associated reflection, and disk emission is sufficient to describe the 0.80.87878 keV spectrum in the SIMS, unlike the HIMS, where two distinct Comptonizing regions were required. The best-fit spectral parameters obtained from M4 are listed in Table 3, and the corresponding spectral data and models are shown in Figure 6.
@@
The Galactic absorption column density, $N_{\rm{H}}$, is found to be $\sim 0.26\times 10^{22}\rm{cm^{-2}}$. The best-fit values of $kT_{\rm{in}}$ and $\Gamma$ indicate that the inner accretion disk has become significantly hotter ($\sim 0.9$ keV), and the continuum has steepened, consistent with the source being in the SIMS. The electron temperature, $kT_{\rm{e}}$, could not be constrained, likely due to the absence of a clear high-energy cutoff. Compared to the HIMS, the inner disk radius has decreased substantially, reaching close to the ISCO with $R_{\rm{in}}\lesssim 2.6~R_{\rm{isco}}$. Additionally, the estimated radius of $\sim 7-8~r_{\rm{g}}$ from the diskbb normalization following the similar procedure adopted in the HIMS also suggests a relatively closer extent of the disk. The disk ionization parameter is also found to be higher than that in the HIMS. As the disk density could not be constrained, it was fixed at its best-fit value of $10^{17}~\rm{cm^{-3}}$, which is slightly higher than the value inferred during the HIMS. These spectral parameters together suggest that the source transitioned to the SIMS during Obs2.
@@
We next replaced the diskbb component with the relativistic disk model kerrbb (Li2005ApJS..157..335L), which describes a geometrically thin, steady-state accretion disk around a Kerr black hole. The inner disk radius was fixed at the ISCO, and a zero-torque boundary condition was assumed at the inner edge. The spectral hardening factor was set to the canonical value of 1.7 (Shimura1995ApJ...445..780S), and disk self-irradiation was included. The black hole spin, $a$, and disk inclination angle, $i$, were tied between the kerrbb and relxillCp components and allowed to vary. Other free parameters of kerrbb are the mass accretion rate ($\dot{M}$), mass of the black hole ($M_{\rm{BH}}$), and distance to the source ($D$). This model, plabs*tbabs(thcomp*kerrbb+relxillCp) (now referred to as M5), provides a better fit compared to the previous model M4, with $\chi^{2}/dof=470.9/519$. The best-fit spectral parameters are given in Table 3, and the corresponding spectral data with the best-fit model is shown in Figure 6.
@@
We obtain the best-fit black hole spin, $a=0.79^{+0.03}_{-0.13}$, mass, $M_{\rm{BH}}=10.45^{+7.7}_{-3.0}~M_{\odot}$, and disk inclination angle, $i=46.6^{+3.2}_{-4.7}$. We also estimated the black hole mass using the mass function proposed by Mata2025A&A...693A.129M, defined as $f\equiv M_{1}\sin^{3}i/(1+M_{2}/M_{1})^{2}=2.77\pm 0.09M_{\odot}$, where $M_{1}$ and $M_{2}$ denote the masses of the black hole and the donor star, respectively. For this, we adopted the value of inclination angle, $i$, from this model M5 and considered $M_{2}$ varying between $0.2–0.78~M_{\odot}$ (Mata2025A&A...693A.129M; zdziarski2025ApJ...986L..35Z). The black hole mass estimated from the mass function lies in the range $\sim 7.8-9.4~M_{\odot}$. In addition, the derived value of the distance to the source obtained from this model is $D=3.47^{+1.32}_{-0.49}$ kpc. Furthermore, the inferred value of the mass accretion rate, $\dot{M}=1.52\pm 0.08$ g/s, corresponds to the luminosity of $\sim 0.1~L_{\rm{Edd}}$, assuming a black hole mass of $M_{\rm{BH}}=10~M_{\odot}$. The other best-fit spectral parameters, including $\Gamma$, $kT_{\rm{e}}$, $\log\xi$, and $\log N$, from the model M5 are well consistent with those estimated from the previous model M4.
@@
In our next model, we used another relativistic disk model bhspec to model the thermal emission from the disk by replacing kerrbb in M5. Like kerrbb, the inner disk is assumed to extend down to the ISCO in bhspec. The primary difference between these two relativistic disk models lies in the treatment of disk emission. While kerrbb considers a color-corrected blackbody approximation, bhspec employs detailed atmospheric-like calculations for each disk annulus to calculate the vertical disk structure and radiative transfer processes self-consistently (Davis2005ApJ...621..372D; Davis2006ApJS..164..530D). bhspec is available as two table models for different values of the viscosity parameter, $\alpha$, of $0.01$ and $0.1$ that consider the black hole spin parameter in the range of 00.990-0.99 and 00.80-0.8, respectively. Since $\alpha=0.1$ is more appropriate for BHXRBs, we adopted the second table model even though it allows a smaller range of black hole spin. The black hole spin ($a$) and disk inclination angle ($i$) were tied between bhspec and relxillCp, and allowed to vary. The other free parameters of bhspec are black hole mass ($M_{\rm{BH}}$), and luminosity of the accretion disk ($L/L_{\rm{Edd}}$). We varied the bhspec normalization, defined as $(10~{\rm{kpc}}/D)^{2}$ (where $D$ is the source distance), over a range corresponding to source distances between 2 and 6 kpc. This model plabs*tbabs(thcomp*bhspec+relxillCp) (referred to as M6) provided a good and statistically acceptable fit with $\chi^{2}/dof=486.8/520$. The best-fit spectral parameters are listed in Table 3, and the spectra, along with the best-fit model, are shown in Figure 6. We find a black hole mass of $M_{\rm{BH}}=19.5^{+1.4}_{-6.0}~M_{\odot}$ and a poorly constrained source distance of $D=6.0_{-0.8}$ kpc. The obtained values of disk luminosity and inclination angle are found to agree with those inferred from M5 within the error bars. Other best-fit spectral parameters also appear to be broadly consistent with those obtained from the models M4 and M5.
Figure 7: Upper panel: PDSs from the NICER, and Insight-HXMT/HE observations, and CPDS from the NuSTAR observation derived in different energy bands during the HIMS. Lower panel: PDSs from the NICER observation derived in two different energy bands during the SIMS. Some of the PDSs are rebinned for plotting purposes only.
@@ 3.2 Timing Analysis
To investigate the source variability, we extracted PDSs across multiple energy bands during both the HIMS and SIMS. In the HIMS, we obtained white-noise subtracted and rms-normalized PDSs in the 0.50.522 keV and 551010 keV bands from NICER data, and in the 30305050 keV, 50508080 keV, and 8080120120 keV bands from HXMT/HE data using the powspec task of the XRONOS package. For HXMT/HE data, the background-subtracted lightcurves were used to derive the PDSs. Additionally, we derived cross power density spectra (CPDSs) in the 10103030 keV band from NuSTAR observations using the HENDRICS package (Bachetti2015ApJ...800..109B), combining the two focal plane modules and subtracting the white noise contribution. The CPDS method involves performing Fourier transforms of two simultaneous light curves from the fpma and fpmb modules and computing their complex cross-spectrum (cospectrum). The real part of this cospectrum serves as a good alternative to a white-noise-subtracted PDS, as it effectively suppresses uncorrelated Poisson noise and properly handles data gaps (see Bachetti2015ApJ...800..109B, for a detailed description). AstroSat/CZTI data were excluded from the timing analysis due to their low signal-to-noise ratio, likely resulting from the relatively small effective area and limited sensitivity of the instrument (Singh2025MNRAS.536.1323S). In the SIMS, PDSs were derived only from NICER data in the 0.50.522 keV and 551010 keV bands. The PDSs and CPDS corresponding to both spectral states are presented in Figure 7.
@@
During the HIMS, the power of the PDS at soft energies is lower across all frequencies, increasing gradually with energy and then saturating beyond $\sim 30$ keV. Notably, the CPDS in the 103010-30 keV band derived from the NuSTAR observation shows relatively lower power than the PDS in the 5105-10 keV band, which could be due to the effect of dead-time in the NuSTAR detectors. Regardless of the energy ranges, a clear presence of a QPO at $\sim 1.3$ Hz accompanied by an upper-harmonic at $\sim 2.7$ Hz is observed in all the PDSs and CPDS in the HIMS. A similar presence of a QPO along with upper-harmonic using AstroSat/LAXPC data in the HIMS of this source was also reported by Nandi2024MNRAS.531.1149N. In the SIMS, the overall shape of the PDSs remains similar. However, the QPO frequency has now shifted to a relatively higher frequency of $\sim 6.6$ Hz, and any prominent signature of the presence of the upper-harmonic up to $10$ Hz was not noticed. The energy-dependent behavior of the PDSs across both spectral states indicates that the low-energy photons, primarily originating from the disk, exhibit significantly lower variability compared to the higher-energy photons, which are dominated by coronal emission.
@@
We also investigated fractional rms variability amplitude as a function of photon energy for the 0.01100.01-10 Hz frequency range using both the observations in the HIMS and SIMS. For the HIMS, we derived the rms-energy spectra over the 0.51200.5-120 keV band using data from NICER, NuSTAR, and HXMT/HE, whereas NICER and NuSTAR data were employed in the SIMS to derive the rms-energy spectrum in the 0.5300.5-30 keV band. We utilized the spectral and timing analysis package Stingray (Huppenkothen2019JOSS....4.1393H; Huppenkothen2019ApJ...881...39H) to derive the rms-energy spectra from NICER data. For the NuSTAR and HXMT/HE data, we first extracted CPDSs and PDSs at different energy bands and then calculated the fractional rms for the 0.01100.01-10 Hz frequency range from each CPDS and PDS. The rms-energy spectra in two distinct spectral states are shown in Figure 8. In both the HIMS and SIMS, the fractional rms variability amplitude appears to be low ($\sim 5\%$) at soft energies, dominated by the thermal disk emission, and gradually increases up to $\sim 30-40\%$ at hard energies, where the coronal emission dominates. We note that the NuSTAR data demonstrate significantly lower variability than the NICER data around the same energy bins, which is likely due to the effect of dead time in the NuSTAR detectors, reducing the variability amplitude artificially (M. Bachetti, private communication; Bachetti2018ApJ...853L..21B; Zdziarski2024ApJ...967L...9Z). It is worth noting that the effects of dead time are significant for bright sources, but become progressively less pronounced for sources of lower intensity. The non-constant and very long ($\sim 2.5$ ms) nature of the NuSTAR dead time further complicates its accurate correction (Bachetti2018ApJ...853L..21B). To further investigate this effect, we used the HXMT/ME data that has a substantial overlap with the NuSTAR data within a certain energy range. We found that the fractional rms derived from the HXMT/ME data is significantly higher than those from the NuSTAR data, and closely matches the measurements from the NICER data. This further confirms that the low rms in the NuSTAR data is due to the effect of the dead time.
4 Discussions

### v2

@@ III.1.1 HIMS
The apparent excess at $\gtrsim 100$ keV further suggests that models with thermal Comptonization and disk reflection alone are insufficient to account for the high-energy emission and may be attributable to a power-law tail in the distribution of hot coronal electrons. The electron distribution is compps hybrid, with the Lorentz factor ($\gamma_{\rm{min}}$) defining the transition point from a thermal to a power-law electron distribution. We allowed $\gamma_{\rm{min}}$ to vary freely while fixing $\gamma_{\rm{max}}=1000$ and keeping the electron power-law index fixed at 2. This model, referred to as M3, which assumes a hybrid electron distribution with Maxwellian and non-thermal above a specific value of $\gamma_{\rm{min}}$, provides a perfect and significantly improved spectral fit in the 0.72000.7-200 keV band with $\chi^{2}$/dof=601.5/681=601.5/681 with $\Delta\chi^{2}\approx-285$ for 1 d.o.f. The best-fit spectral parameters obtained from M3 are listed in Table 2, and the best-fit spectral data with the model components are shown in Figure 4. We find $\gamma_{\rm{min}}$ to be $\sim 1.33$, implying a significant contribution from the non-thermal distribution. The other best-fit spectral parameters obtained with model M3 are in close agreement with those from our earlier models, assuming a purely thermal electron distribution up to $78$ keV.
III.1.2 SIMS
@@
Thcomp $\Gamma$ $2.36\pm 0.03$ $2.33^{+0.10}_{-0.06}$ $2.35^{+0.03}_{-0.04}$
$kT_{\rm{e}}$ (keV) $998.2^{p}_{-500.4}$ $668.5^{p}_{-554.6}$ $995^{p}_{-502}$
$C_{\rm{f}}$ $0.25^{+0.04}_{-0.06}$ $0.2^{+0.2}_{-0.1}$ $0.23^{+0.05}_{-0.06}$
Kerrbb $a$  $\cdots$ $0.79^{+0.07}_{-0.15}$  $\cdots$
$i\arcdeg$  $\cdots$ $46.6^{+6.7}_{-9.2}$  $\cdots$
$\dot{M}_{\rm d}$ ($\times 10^{18}~\rm{g/s}$)  $\cdots$ $1.5\pm 0.2$  $\cdots$
$M_{\rm{BH}}$ ($M_{\odot}$)  $\cdots$ $10.3^{+5.5}_{-2.5}$  $\cdots$
$D$ (kpc)  $\cdots$ $3.4^{f}$  $\cdots$
BHSPEC $\alpha=0.1$] $M_{\rm{BH}}$ ($M_{\odot}$)  $\cdots$  $\cdots$ $10.2^{+1.3}_{-0.4}$
$L/L_{\rm{Edd}}$  $\cdots$  $\cdots$ $0.12^{+0.02}_{-0.01}$
$a$  $\cdots$  $\cdots$ $0.80^{p}_{-0.03}$
$i\arcdeg$  $\cdots$  $\cdots$ $47.2\pm 2.3$
$D$ (kpc)  $\cdots$  $\cdots$ $3.4^{f}$
RelxillCp (refl. only) $a$ $0.98^{p}_{-0.29}$  $\cdots$  $\cdots$
$i\arcdeg$ $40^{f}$  $\cdots$  $\cdots$
$R_{\rm{in}}$ ($r_{\mathrm{isco}}$) $1.7^{+0.9}_{p}$ $1^{f}$ $1^{f}$
$q$ $3.7^{+0.6}_{-0.4}$ $6.6^{p}_{-3.0}$ $6.0^{+2.1}_{-1.2}$
$\log\xi$ $4.09^{+0.04}_{-0.11}$ $4.26^{+0.12}_{-0.21}$ $4.20^{+0.04}_{-0.10}$
@@
$K_{\rm{NICER}}$ $0.82\pm 0.02$ $0.81^{+0.01}_{-0.02}$ $0.81\pm 0.02$
$\Delta\Gamma_{\rm{NuSTAR~B}}$ $002^{+0.006}_{p} 0.0$$03^{+0.006}_{p} 0.0$$002^{+0.006}_{p}$
@@
  • Note. M4: plabs*tbabs(thcomp*diskbb+mbknpo*relxillCp), M5: plabs*tbabs(thcomp*kerrbb+mbknpo*relxillCp), M6: plabs*tbabs(thcomp*bhspec+mbknpo*relxillCp). f – indicates fixed parameters. $p$ – indicates that the parameter is pegged at the model-defined highest or lowest bound during error estimation.
@@
We then employed the more physically motivated and relativistically blurred reflection model, relxillCp, to perform broadband reflection spectroscopy, replacing the ireflect and gaussian components. Here, relxillCp is considered only as a reflection component by fixing the reflection fraction parameter at 1-1. Common parameters, such as $\Gamma$ and $kT_{\rm{e}}$, were tied between the relxillCp and thcomp components. The inclination angle and the iron abundance were also kept fixed at $40\arcdeg$ and solar, respectively. Other free spectral parameters of the relxillCp model are the emissivity index ($q$), log of disk ionization ($\log\xi$), log of disk density ($\log\xi$), and normalization. Additionally, we varied the black hole spin ($a$) and the inner disk radius ($R_{\rm{in}}$) independently, accounting for the SIMS. The resulting model plabs*tbabs(thcomp*diskbb+relxillCp) (now referred to as M4) provides a good and statistically acceptable fit with $\chi^{2}/dof=506.5/521$. This further suggests that a model consisting of a single Comptonizing region, its associated reflection, and disk emission is sufficient to describe the 0.80.87878 keV spectrum in the SIMS, unlike the HIMS, where two distinct Comptonizing regions were required. The best-fit spectral parameters obtained from M4 are listed in Table 3, and the corresponding spectral data and models are shown in Figure 6.
@@
The Galactic absorption column density, $N_{\rm{H}}$, is found to be $\sim 0.26\times 10^{22}\rm{cm^{-2}}$. The best-fit values of $kT_{\rm{in}}$ and $\Gamma$ indicate that the inner accretion disk has become significantly hotter ($\sim 0.9$ keV), and the continuum has steepened, consistent with the source being in the SIMS. The electron temperature, $kT_{\rm{e}}$, could not be constrained, likely due to the absence of a clear high-energy cutoff. Compared to the HIMS, the inner disk radius has decreased substantially, reaching close to the ISCO with $R_{\rm{in}}\lesssim 2.6~R_{\rm{isco}}$. Additionally, the estimated radius of $\sim 7-8~r_{\rm{g}}$ from the diskbb normalization following the similar procedure adopted in the HIMS also suggests a relatively closer extent of the disk. The disk ionization parameter is also found to be higher than that in the HIMS. Because disk density could not be constrained, it was fixed at its best-fit value of $10^{17}~\rm{cm^{-3}}$, which is slightly higher than the value inferred during HIMS. These spectral parameters together suggest that the source transitioned to the SIMS during Obs2.
@@
We next replaced the diskbb component with the relativistic disk model kerrbb (Li et al., 2005), which describes a geometrically thin, steady-state accretion disk around a Kerr black hole. The inner disk radius was fixed at the ISCO, and a zero-torque boundary condition was assumed at the inner edge. The spectral hardening factor was set to the canonical value of 1.7 (Shimura and Takahara, 1995), and disk self-irradiation was included. The black hole spin, $a$, and disk inclination angle, $i$, were tied between the kerrbb and relxillCp components and allowed to vary. Other free parameters of kerrbb are the disk mass accretion rate ($\dot{M}_{\rm d}$), and mass of the black hole ($M_{\rm{BH}}$).
We note that the kerrbb model is based on the thin-disk approximation; therefore, the fitted spectrum is insensitive to the Eddington ratio (which, in reality, would affect the disk scale height). For a given spin, this approximation yields spectra that depend only on the disk’s maximum temperature, $T$, and flux, $F$. Consequently, we have $T\propto\dot{M}_{\rm d}/M_{\rm BH}^{2}$ and $F\propto\dot{M}_{\rm d}/D^{2}$. This implies that changes in $M_{\rm BH}$, $D$, and $\dot{M}_{\rm d}$ do not affect the fitted spectrum as long as the scaling of $M_{\rm BH}\propto D\propto\dot{M}_{\rm d}^{1/2}$ holds. Thus, to break this degeneracy, we fixed the distance at the best-fit value of Mata Sánchez et al. (2025), $D=3.4$ kpc (see also discussion in Zdziarski et al. 2025). This model, plabs*tbabs(thcomp*kerrbb+relxillCp) (now referred to as M5), provides a better fit compared to the previous model M4, with $\chi^{2}/dof=470.8/520$. The best-fit spectral parameters are given in Table 3, and the corresponding spectral data with the best-fit model are shown in Figure 6.
@@
We obtain the best-fit black hole spin, $a=0.79^{+0.07}_{-0.15}$, mass, $M_{\rm{BH}}=10.3^{+5.5}_{-2.5}~M_{\odot}$, and disk inclination angle, $i=46.6^{+6.7}_{-9.2}\arcdeg$. We also estimated the black hole mass using the mass function proposed by Mata Sánchez et al. (2025), defined as $f\equiv M_{\rm{BH}}\sin^{3}i/(1+M_{2}/M_{1})^{2}=2.77\pm 0.09M_{\odot}$, where $M_{2}$ denotes the mass of the donor star. For this, we adopted the value of inclination angle, $i$, from this model M5 and considered $M_{2}$ varying between $0.2–0.78~M_{\odot}$ (Mata Sánchez et al., 2025; Zdziarski et al., 2025). The black hole mass estimated from the mass function lies in the range $\approx 7$$12~M_{\odot}$. Furthermore, the fitted values of $M$ and the disk accretion rate, $\dot{M}_{\rm d}=1.5\pm 0.2\times 10^{18}$ g/s, correspond to the disk luminosity of $L_{\rm disk}\approx 0.13L_{\rm{Edd}}$, assuming the disk accretion efficiency corresponding to the best-fit value of the spin (Cunningham, 1975). We note here the scalings of $M_{\rm{BH}}\propto D$, $\dot{M}_{\rm d}\propto D^{2}$ and $L/L_{\rm Edd}\propto D$. The other best-fit spectral parameters, including $\Gamma$, $kT_{\rm e}$, $\log\xi$, and $\log N$, from the model M5 are well consistent with those estimated from the previous model M4.
@@
In our next model, M6, we used the relativistic disk model bhspec to represent the disk’s thermal emission, replacing kerrbb in M5. Like kerrbb, the inner disk is assumed to extend down to the ISCO in bhspec. The primary difference between these two relativistic disk models lies in how they treat disk emission. While kerrbb considers a color-corrected blackbody approximation, bhspec employs detailed atmospheric-like calculations for each disk annulus to calculate the vertical disk structure and radiative transfer processes self-consistently (Davis et al., 2005; Davis and Hubeny, 2006). bhspec is available in two table models, with the viscosity parameters $\alpha=0.01$ and $0.1$ and covering the black hole spin parameters 00.990-0.99 and 00.80-0.8, respectively. Since $\alpha=0.1$ is more appropriate for BHXRBs, we adopted the second table model even though it allows a smaller range of black hole spin. The black hole spin ($a$) and disk inclination angle ($i$) were tied between bhspec and relxillCp, and allowed to vary. The other free parameters of bhspec are black hole mass ($M_{\rm{BH}}$), and luminosity of the accretion disk ($L/L_{\rm{Edd}}$). We fixed the bhspec normalization, defined as $(10~{\rm{kpc}}/D)^{2}$, to a value corresponding to $D=3.4$ kpc (Mata Sánchez et al., 2025; Zdziarski et al., 2025). Allowing the distance to vary resulted in a relatively larger value ($D\gtrsim 5.2$ kpc) and, consequently, a higher black hole mass. This model, plabs*tbabs(thcomp*bhspec+relxillCp), provided a good and statistically acceptable fit with $\chi^{2}/dof=489/521$. The best-fit spectral parameters are listed in Table 3, and the spectra, along with the best-fit model, are shown in Figure 6. We find a black hole mass of $M_{\rm{BH}}=10.2^{+1.3}_{-0.4}~M_{\odot}$. The obtained values of disk luminosity and inclination angle are consistent with those inferred from M5 within the error bars. Other best-fit spectral parameters are also broadly consistent with those obtained from models M4 and M5.
Figure 7: Upper panel: PDSs from the NICER and Insight-HXMT/HE observations, and CPDS from the NuSTAR observation derived in different energy bands during the HIMS. Lower panel: PDSs from the NICER observation derived in two different energy bands during the SIMS. Some of the PDSs are rebinned for plotting purposes only.
@@ III.2 Timing Analysis
To investigate the source variability, we extracted PDSs across multiple energy bands during both the HIMS and SIMS. In the HIMS, we obtained white-noise subtracted and rms-normalized PDSs in the 0.50.522 keV and 55–-$10$ keV bands from NICER data, and in the 30305050 keV, 50508080 keV, and 8080–-$120$ keV bands from HXMT/HE data using the powspec task of the XRONOS package. For HXMT/HE data, the background-subtracted lightcurves were used to derive the PDSs. Additionally, we derived cross power density spectra (CPDSs) in the 1010–-$30$ keV band from NuSTAR observations using the HENDRICS package (Bachetti et al., 2015), combining the two focal plane modules and subtracting the white noise contribution. The CPDS method involves performing Fourier transforms of two simultaneous light curves from the fpma and fpmb modules and computing their complex cross-spectrum (cospectrum). The real part of this cospectrum serves as a good alternative to a white-noise-subtracted PDS, as it effectively suppresses uncorrelated Poisson noise and correctly handles data gaps (see Bachetti et al., 2015, for a detailed description). AstroSat/CZTI data were excluded from the timing analysis due to their low signal-to-noise ratio, likely resulting from the instrument’s relatively small effective area and limited sensitivity (Singh et al., 2025). In the SIMS, PDSs were derived only from NICER data in the 0.50.522 keV and 551010 keV bands. The PDSs and CPDS corresponding to both spectral states are presented in Figure 7.
@@
During the HIMS, the power of the PDS at soft energies is lower across all frequencies, increasing gradually with energy and then saturating beyond $\sim 30$ keV. Notably, the CPDS in the 103010-30 keV band derived from the NuSTAR observation shows relatively lower power than the PDS in the 5–10 keV band, which could be due to the effect of dead-time in the NuSTAR detectors. Regardless of the energy ranges, a clear presence of a QPO at $\sim 1.3$ Hz accompanied by an upper-harmonic at $\sim 2.7$ Hz is observed in all the PDSs and CPDS in the HIMS. A similar presence of a QPO, along with an upper harmonic, in the HIMS of this source using AstroSat/LAXPC data was also reported by Nandi et al. (2024). In the SIMS, the overall shape of the PDSs remains similar. However, the QPO frequency has now shifted to a relatively higher frequency of $\sim 6.6$ Hz, and any prominent signature of the presence of the upper-harmonic up to $10$ Hz was not noticed. The energy-dependent behavior of the PDSs across both spectral states indicates that low-energy photons, primarily originating from the disk, exhibit significantly lower variability than higher-energy photons, which are dominated by coronal emission.
@@
We also investigated the fractional rms variability amplitude as a function of photon energy in the 0.01–10 Hz frequency range, using both HIMS and SIMS observations. For the HIMS, we derived the rms-energy spectra over the 0.5–120 keV band using data from NICER, NuSTAR, and HXMT/HE, whereas NICER and NuSTAR data were employed in the SIMS to derive the rms-energy spectrum in the 0.5300.5-30 keV band. We utilized the spectral and timing analysis package Stingray (Huppenkothen et al., 2019b, a) to derive the rms-energy spectra from NICER data. For the NuSTAR and HXMT/HE data, we first extracted CPDSs and PDSs at different energy bands and then calculated the fractional rms for the 0.01100.01-10 Hz frequency range from each CPDS and PDS. The rms-energy spectra in two distinct spectral states are shown in Figure 8. In both the HIMS and SIMS, the fractional rms variability amplitude appears to be low ($\sim 5\%$) at soft energies, dominated by the thermal disk emission. It gradually increases up to $\sim 30-40\%$ at hard energies, where the coronal emission dominates. We note that the NuSTAR data demonstrate significantly lower variability than the NICER data in the same energy bins, likely due to dead time in the NuSTAR detectors, which artificially reduces the variability amplitude (M. Bachetti, private communication; Bachetti and Huppenkothen, 2018; Zdziarski et al., 2024b). It is worth noting that the effects of dead time are significant for bright sources but become progressively less pronounced for lower-intensity sources. The non-constant and very long ($\sim 2.5$ ms) nature of the NuSTAR dead time further complicates its accurate correction (Bachetti and Huppenkothen, 2018). To further investigate this effect, we used HXMT/ME data that substantially overlap with the NuSTAR data within a specific energy range. We found that the fractional rms derived from the HXMT/ME data is significantly higher than that from the NuSTAR data, and closely matches the measurements from the NICER data. This further confirms that the low rms in the NuSTAR data is due to dead time.
IV Discussions